differentiate the following function with respect to x
(1) sin^-1(sqrt(1-(x)^2)),0
(2) tan^-1(x/sqrt(a^2 – x^2)), -a

Grade:12

1 Answers

Rinkoo Gupta
9 years ago

y=sin-1v(1-x2)



dy/dx=1/v1-(1-x2) .{-2x/v(1-x2)}



=-1/v(1-x2)

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty